OpenStudy (anonymous):

convert the following equation from polar to rectangular form: theta=2pi/3

OpenStudy (dumbcow):

use \[\tan \theta = \frac{y}{x}\]

OpenStudy (jdoe0001):

\(\bf \theta=\cfrac{{\color{red}{ 2\pi}}}{{\color{blue}{ 3}}}\qquad tan(\theta)=\cfrac{{\color{red}{ y}}}{{\color{blue}{ x}}}\qquad rectangular\implies (\square ,\square )\)

OpenStudy (anonymous):

I know that the answer is: \[\sqrt{3}x + y = 0\] but unsure the step to get there.

OpenStudy (dumbcow):

sub theta into equation tan theta = y/x what is tan(2pi/3) ?

OpenStudy (anonymous):

Aha. I think I understand now. Thanks.

OpenStudy (dumbcow):

yw :)